OpenStudy (anonymous):

the area of a rectangle is 99yd^2 and the length is 7yd then double the widtth

OpenStudy (anonymous):

question is real unclear what does "the length is 7 yds then double the width" mean?

OpenStudy (anonymous):

can you post a screenshot or something? hard to guess what is means

OpenStudy (anonymous):

y yards more than twice the width?

OpenStudy (anonymous):

thats all it says thats why it confused me. I dont understand what I am supposed to be doing.

OpenStudy (anonymous):

can you post a picture? or write it again verbatim? the question is really unclear

OpenStudy (anonymous):

my guess: the area of a rectangle is 99 square yards the length is 7 more than twice the width find the dimensions but that is just a guess

OpenStudy (anonymous):

omg im an idiot your right. I didnt read the whole thing correctly. I hate online clases

OpenStudy (anonymous):

ok now we are getting somewhere lets call the width \(x\) so the length is \(2x+7\)

OpenStudy (anonymous):

since area is length times width you have \[x(2x+7)=99\] solve for \(x\)

OpenStudy (anonymous):

i.e. \[2x^2+7x-99=0\] if you are lucky it factors

OpenStudy (anonymous):

is it 11??

OpenStudy (anonymous):

man my algebra stinks

OpenStudy (anonymous):

it should be \[x(2x+7)=99\ 2x^2+14x=--\\ 2x^2+14x-99=0\]

OpenStudy (anonymous):

mine too lol I havent been in school in 8 years so I had to take college algebra

OpenStudy (anonymous):

no that is not good either

OpenStudy (anonymous):

if it is an on line class can you post a screen shot?

OpenStudy (anonymous):

how do i screen shot my laptop?

OpenStudy (anonymous):